OpenStudy (anonymous):

Evaluate the function for the given input values. For f(x) = 3x -1, when x = -2 and when x=7. Show both answers.

OpenStudy (anonymous):

you just replace the x with its value meaning f(-2)= 3(-2)-1= -2f=(-6)-1= -2f=-6= f/-2=0.....-6/-2= 3 f=3

OpenStudy (anonymous):

f(x)=3x-1 f(7)=3(7)-1 7f=21-1 7f=20 f/7=0..... 20/7=2.85 f=2.85